OpenStudy (anonymous):

What is the term for a relation in which each input value corresponds to exactly one output value?

Directrix (directrix):

Click here for a definition that will help answer the question: http://tinyurl.com/kuftetb @thannoi123

OpenStudy (anonymous):

f(x) = x/2 ("f of x is x divided by 2") is a function, because for every value of "x" you get another value "x/2", so: * f(2) = 1 * f(16) = 8 * f(-10) = -5 ?

OpenStudy (anonymous):

HW

OpenStudy (anonymous):

can u help me?

Directrix (directrix):

The first question you posted: what did you get for an answer? What is the term for a relation in which each input value corresponds to exactly one output value?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):